Suppose a population contains 100,000 individuals, of which 500 are carriers of a certain genetic defect.
Which, if any, of the following is NOT an appropriate way to express this result as a portion of the total? Why?!
A- One half of one percent (.5%)of all individuals in this population are carriers.
B-The proportion in this population who are carriers is .005
C- The probability that a person randomly selected from this population carrier is .5%
D- All of the above are appropriate.

Answer by stanbon(75887)

Suppose a population contains 100,000 individuals, of which 500 are carriers of a certain genetic defect.
----
portion = 500/100000 = 5/1000 = 0.005 = 0.5%
------------------------------------------------------
Which, if any, of the following is NOT an appropriate way to express this result as a portion of the total? Why?!
A- One half of one percent (.5%)of all individuals in this population are carriers.
B-The proportion in this population who are carriers is .005
C- The probability that a person randomly selected from this population carrier is .5%
D- All of the above are appropriate.
----------
Answer: D
===============
Cheers,
Stan H.
============